# Annealed Multiple Choice Learning: Overcoming limitations of Winner-takes-all with annealing

### v1

@@ 5.1 UCI datasets
Baselines. In our result tables, we also include data from three baseline methods detailed in Table 1 of Lakshminarayanan et al.’s paper [37], which serves as a reference. These baselines include Probabilistic Back Propagation [28] (denoted ‘PBP’), and Monte Carlo Dropout [20] (denoted ‘MC-dropout’). As additional baselines, we include the standard score-based MCL (e.g., [41]). We also include the Relaxed-MCL variant [60] with $\varepsilon=0.5$. Our method (aMCL), was trained with $T_{\mathrm{ini}}=0.5$ and a linear scheduler. Both aMCL and Relaxed-MCL were trained for 1,000 epochs. Each MCL system was trained with $n=5$ hypotheses.
Metrics. We computed the following metrics on the UCI datasets. Let $d$ denote the squared Euclidean distance: $d(\hat{y_{i}},y_{i})=\lVert y_{i}-\hat{y_{i}}\rVert^{2}$, and $N$ the number of samples in each dataset. The RMSE ()($\downarrow$)( ↓ ) is defined as $\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i}d(\hat{y}_{i},y_{i})}$, where $\hat{y}_{i}$ denotes the estimated conditional mean. The latter was computed with $\sum_{k=1}^{n}\gamma_{k}(x)f_{k}(x)$ for the MCL variants. The results of this experiment are summed up in Table 1 (RMSE), and also in Table 3 (Distortion) of the appendix.
Comparison of aMCL and MCL. We can observe that aMCL performs comparably or outperforms vanilla MCL in most settings, both in terms of distortion and RMSE. This outcome supports the claims made in the paper and is especially promising, given that the temperature scheduler was not specifically optimized for each dataset.
Comparison of aMCL and standard quantization baselines. We observe that aMCL performs comparably to, and occasionally surpasses the standard baselines on the RMSE metric. This is worth noticing, given that optimizing RMSE is not the primary objective of aMCL training. For instance, the RMSE performance of MCL is significantly worse in most datasets.
Comparison with Relaxed-MCL. Finally, we compare aMCL and Relaxed-MCL, since aMCL can be interpreted as an adaptative version of Relaxed-MCL. First, our observations suggest that aMCL is typically on par with, and occasionally surpasses Relaxed-MCL in terms of distortion across various settings. Second, we notice the very strong performance of Relaxed-MCL in terms of RMSE. We believe that this is due to the bias of Relaxed-MCL toward the conditional barycenter of the target distribution, and the fact that RMSE evaluates the quality of this barycenter estimation.
These results strongly support the use of aMCL as a quantization algorithm in practical settings. To further evaluate its effectiveness, we also apply aMCL to a realistic task, namely speech separation.
Table 1: Results on UCI regression benchmark datasets comparing RMSE. Best results are in bold, second bests are underlined.
@@
Datasets PBP MC Dropout Deep Ensembles Relaxed-MCL MCL aMCL
Boston 3.01 $\pm$ 0.18 2.97 $\pm$ 0.85 3.28 $\pm$ 1.00 2.85 $\pm$ 0.57 3.54 $\pm$ 1.16 3.12 $\pm$ 0.68
Concrete 5.67 $\pm$ 0.09 5.23 $\pm$ 0.53 6.03 $\pm$ 0.58 4.91 $\pm$ 0.65 6.02 $\pm$ 0.65 4.99 $\pm$ 0.63
Energy 1.80 $\pm$ 0.05 1.66 $\pm$ 0.19 2.09 $\pm$ 0.29 1.03 $\pm$ 0.21 2.53 $\pm$ 0.99 0.80 $\pm$ 0.33
Kin8nm 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00 0.09 $\pm$ 0.00 0.08 $\pm$ 0.00 0.10 $\pm$ 0.01 0.08 $\pm$ 0.00
Naval 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Power 4.12 $\pm$ 0.03 4.02 $\pm$ 0.18 4.11 $\pm$ 0.17 3.99 $\pm$ 0.16 4.18 $\pm$ 0.16 4.07 $\pm$ 0.16
Protein 4.73 $\pm$ 0.01 4.36 $\pm$ 0.04 4.71 $\pm$ 0.06 4.20 $\pm$ 0.02 4.39 $\pm$ 0.10 4.26 $\pm$ 0.04
Wine 0.64 $\pm$ 0.01 0.62 $\pm$ 0.04 0.64 $\pm$ 0.04 0.63 $\pm$ 0.04 0.63 $\pm$ 0.04 0.67 $\pm$ 0.05
Yacht 1.02 $\pm$ 0.05 1.11 $\pm$ 0.38 1.58 $\pm$ 0.48 1.17 $\pm$ 0.32 3.28 $\pm$ 1.39 0.85 $\pm$ 0.25
Year 8.88 $\pm$ NA 8.85 $\pm$ NA 8.89 $\pm$ NA 8.91 $\pm$ NA 9.09 $\pm$ NA 9.20 $\pm$ NA
@@ 5.2 Application to speech separation
Speech separation consists of isolating each speaker’s signal from a mixture in which they are simultaneously active. This task is of major interest for automatic speech processing applications [49, 43, 67]. In these experiments, we explore the application of MCL and the proposed aMCL to the source separation task. An extensive description of the experiments is proposed in Appendix D.2.
@@ 5.2.1 Experimental setting
Dataset. Source separation experiments are conducted on the Wall Street Journal dataset [29], the current standard benchmark for speech separation. We focus on the 2- and 3- speaker mixture scenarios, with each scenario including 20,000 training, 5,000 validation, and 3,000 testing mixtures.
Model architecture. The source separation task is solved using the Dual-Path Recurrent Neural Network (DPRNN) architecture [45]. DPRNN has been extensively used in speech separation and provides a favorable balance between performance and number of trainable parameters.
Separation metrics. We use the Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) to measure the separation quality [66, 38]. There is an ambiguity in finding the best assignment between predicted and active sources. The PIT SI-SDR loss [36] initially addresses this issue. We propose to use MCL and our new variant aMCL to perform this matching (See Appendix D).
@@ 5.2.2 Results
Table 2: PIT SI-SDR performance ($\uparrow$) on source separation in the 2- and 3- speaker scenarios.
@@
Comparing PIT with MCL and aMCL. Table 2 presents the average source separation results in the 2- and 3-speaker scenarios. First, aMCL demonstrates performance equivalent or better to that of MCL. Both models can be used for the separation task. However, we observed that MCL is subject to hypothesis collapse for some training seeds, while aMCL is more robust to initialization. Second, aMCL perform equivalently to PIT in both 2- and 3-speaker scenarios. Moreover, in the MCL and aMCL settings, we can let the number of predictions $n$ exceed the number of sources $m$. This could be leveraged to improve separation metrics (cf. Appendix D.3.2). Finally, aMCL improves the algorithmic complexity of PIT from $\mathcal{O}(m^{3})$ to $\mathcal{O}(mn)$ (cf. Appendix D.3.1). These results make aMCL stand as a good alternative to PIT.
Observing phase transition. In the context of the Euclidean distance metric, the theoretical analysis 4.3 and synthetic experiments (see Figure 3) have highlighted a phenomenon of phase transition for aMCL. Here, we analyze the validation loss trajectory as a function of the temperature for different initial temperatures, in the case of an exponential scheduler. The curves with the two higher initial temperatures in Figure 6 exhibit a plateau until a given temperature. After this critical point, the loss decreases. Although the SI-SDR is non-Euclidean, this behavior resembles that observed for Euclidean metrics. This is detailed in Section D.3.3 of the Appendix.
@@ 6 Conclusion
This article introduces aMCL, a novel training method that combines deterministic annealing and the Winner-Takes-all training scheme to address two key issues of MCL: hypothesis collapse and convergence toward a suboptimal local minimum of its quantization objective. We provide a detailed analysis of aMCL’s training dynamics. Moreover, drawing on statistical physics and information theory, we provide insights into the trajectory of the aMCL predictions during training. In particular, we exhibit a phase transition phenomenon and formally establish its connection to the rate-distortion curve. We validate our analysis with experiments on synthetic data, on the UCI datasets, and on a real-world speech separation task. This demonstrates that aMCL is a theoretically grounded alternative to MCL in diverse settings. Future work includes a detailed analysis of the temperature schedule’s impact, the derivation of performance bounds, and a thorough examination of our algorithm’s convergence, particularly at finite temperature.
Limitations. First, aMCL introduces a temperature schedule: this is a challenging hyperparameter tightly linked to the optimizer and its learning rate. The derivation of optimal schedules is left to future work. Second, annealing requires a slow temperature schedule to maintain model performance, leading potentially to longer training times.
@@ A.1 Notations and motivations
Following the main paper notations, let $\mathcal{X}\subset\mathbb{R}^{d_{1}}$ and $\mathcal{Y}\subset\mathbb{R}^{d_{2}}$ denote the input and target spaces respectively. We note $p(x,y)$ the density of the joint data distribution on $\mathcal{X}\times\ \mathcal{Y}$. Let $f=(f_{1},\dots,f_{n})\in\mathcal{F}_{\Theta}\subset\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})$ denote the hypothesis models, which are families of $n$ neural networks with parameters in $\Theta$. Likewise, let $\gamma=(\gamma_{1},\dots,\gamma_{n})\in\mathcal{F}_{\Theta^{\prime}}\subset\mathcal{F}(\mathcal{X},[0,1]^{n})$ denote the score models, with parameters in $\Theta^{\prime}$. We will sometimes write $\mathcal{F}=\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})$.
We denote by $\Delta_{n}$ the set of discrete distributions on $n$ items conditioned by points on $\mathcal{X}\times\mathcal{Y}$ (typically, distributions on the positions $f_{k}$). We denote by $H(q)$ the entropy of a distribution $q$. If $q\in\Delta_{n}$, we write $H(q)=-\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}q(f_{k}|x,y)\log q(f_{k}|x,y)p(x,y)\mathrm{d}x\mathrm{d}y$. If $Z$ is a random variable with density $p(z)$ and $Y$ is another random variable, we define the differential entropy of $Z$ as $H(Z)=-\mathbb{E}_{Z}[\mathrm{log}(p(Z))]$ and its conditional entropy as $H(Z\mid Y)=-\mathbb{E}_{(Z,Y)}[\mathrm{log}(p(Z\mid Y))]$. Mutual information $I(Z,Y)$ between two random variables $Z$ and $Y$ is defined as $I(Z;Y)=H(Z)-H(Z\mid Y)$.
Let $t\in\mathbb{N}$ denote a training step, and $t_{\mathrm{epoch}}\in\mathbb{N}\cup\{\infty\}$ denote the number of training epochs. Let $T:t\mapsto T(t)\geq 0$ be a temperature schedule that decreases with the training step and verifies $\mathrm{lim}_{t\rightarrow t_{\mathrm{epoch}}}T(t)=0$. Note that we consider the temperature to be constant during each epoch. Let $\ell:\mathcal{Y}^{2}\rightarrow\mathbb{R}_{+}$ be the underlying loss function. We will restrict our analysis to the Euclidean squared distance $\ell(\hat{y},y)=\lVert y-\hat{y}\rVert^{2}$ unless otherwise stated. For any vector $z$, we note $z^{t}$ as its vector transposition. We will assume that $\mathcal{Y}$ is bounded. In particular, this ensures that all integrals of the form $\mathbb{E}_{Z\sim p(x,y)}[\phi(Z)]$ are well defined when $\phi$ is a $\mathcal{Y}$-measurable function. Gradient descent involves a learning rate schedule, that we note $\lambda_{t}$.
@@
which can be seen as a generalization of the conditional distortion [53]. The two definitions coincide for fixed distributions, i.e., when $p(y\mid x)$ is independent of $x$. Minimizing (16) is NP-hard [1, 10, 47]. The K-Means algorithm, which is the standard approach for this task, can take an exponential number of steps before converging [65], and is only guaranteed to find a local minimum in general settings [13, 7]. The same applies to WTA: due to its greedy nature, it can be trapped, at any point of its training trajectory, into a local minimum that it will never manage to escape. One common issue is the hypothesis collapse [54, 17], a situation in which some of the hypotheses are left unused, therefore leading to a suboptimal configuration (see Figure 1).

### v3

@@ 5.1.1 Setup
Baselines. In our result tables, we also include data from three baseline methods detailed in Table 1 of Lakshminarayanan et al.’s paper [38] (‘Deep Ensembles’), which serves as a reference. These baselines include Probabilistic Back Propagation [29] (denoted ‘PBP’), and Monte Carlo Dropout [21] (denoted ‘MC-dropout’). As additional baselines, we include the standard score-based MCL (e.g., [42]). We also include the Relaxed-MCL variant [63] with $\varepsilon=0.1$. The impact of $\varepsilon$ is discussed in Appendix D.2. Our method (aMCL), was trained with an exponential scheduler of the form $T(t)=T_{0}\rho^{t}$, with $\rho=0.95$ and $T_{0}=0.5$. Comparison with a linear scheduler is also provided in Appendix D.1. Both aMCL and Relaxed-MCL were trained for 1,000 epochs. Each MCL system was trained with $n=5$ hypotheses.
Metrics. We computed the following metrics on the UCI datasets. Let $d$ denote the squared Euclidean distance: $d(\hat{y}_{i},y_{i})=\lVert y_{i}-\hat{y_{i}}\rVert^{2}$, and $N$ the number of samples in each dataset. The RMSE ()($\downarrow$)( ↓ ) is defined as $\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i}d(\hat{y}_{i},y_{i})}$, where $\hat{y}_{i}$ denotes the estimated conditional mean. The latter was computed with $\sum_{k=1}^{n}\gamma_{k}(x_{i})f_{k}(x_{i})$ for the MCL variants. The results of this experiment are summed up in Table 1 (Distortion), and also in Table 2 (RMSE). Rows are ordered by dataset size $N$, with the intensity of the grey color proportional to $N$ (excluding the Year dataset).
5.1.2 Results
Table 1: Results on UCI regression benchmark datasets comparing Distortion. Experimental setup is described in Section 5.1.1. Relaxed-WTA results were computed with $\varepsilon=0.1$ which strikes a good tradeoff between RMSE and Distortion (see Table 5 in Appendix). The rows are ordered by dataset size $N$. Best results are in bold, second bests are underlined.
Distortion ($\downarrow$)
Datasets Relaxed-WTA ($\varepsilon=0.1$) MCL aMCL $N$
Year 9.09 $\pm$ NA 4.82 $\pm$ NA 4.46 $\pm$ NA 515345
Protein 1.67 $\pm$ 0.16 0.80 $\pm$ 0.02 0.77 $\pm$ 0.03 45730
Naval 4.21e-7 $\pm$ 2.36e-7 1.84e-6 $\pm$ 2.42e-6 5.37e-7 $\pm$ 3.83e-7 11934
Power 2.95 $\pm$ 0.91 2.31 $\pm$ 0.49 2.18 $\pm$ 0.64 9568
Kin8nm 9.32e-4 $\pm$ 7.97e-5 1.00e-3 $\pm$ 1.47e-4 6.81e-4 $\pm$ 8.14e-5 8192
Wine 0.06 $\pm$ 0.02 0.02 $\pm$ 0.01 0.03 $\pm$ 0.01 1599
Concrete 6.91 $\pm$ 2.81 5.13 $\pm$ 1.23 5.71 $\pm$ 1.72 1030
Energy 0.30 $\pm$ 0.12 1.25 $\pm$ 1.25 0.28 $\pm$ 0.09 768
Boston 3.32 $\pm$ 2.84 2.14 $\pm$ 0.49 2.69 $\pm$ 1.39 506
Yacht 1.34 $\pm$ 0.93 3.09 $\pm$ 2.41 1.15 $\pm$ 0.97 308
Table 2: Results on UCI regression benchmark datasets comparing RMSE. Best results are in bold, second bests are underlined. corresponds to reported results from [38].
@@ 5.1.2 Results
Datasets PBP MC Dropout Deep Ensembles Relaxed-WTA ($\varepsilon=0.1$) MCL aMCL $N$
Year 8.88 $\pm$ NA 8.85 $\pm$ NA 8.89 $\pm$ NA 8.97 $\pm$ NA 9.09 $\pm$ NA 9.08 $\pm$ NA 515345
Protein 4.73 $\pm$ 0.01 4.36 $\pm$ 0.04 4.71 $\pm$ 0.06 4.38 $\pm$ 0.02 4.39 $\pm$ 0.10 4.25 $\pm$ 0.02 45730
Naval 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 1.80e-3 $\pm$ 5.66e-4 2.08e-3 $\pm$ 1.18e-3 8.00e-4 $\pm$ 4.04e-4 11934
Power 4.12 $\pm$ 0.03 4.02 $\pm$ 0.18 4.11 $\pm$ 0.17 4.02 $\pm$ 0.18 4.18 $\pm$ 0.16 4.08 $\pm$ 0.20 9568
Kin8nm 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00 0.09 $\pm$ 0.00 0.08 $\pm$ 0.00 0.10 $\pm$ 0.01 0.08 $\pm$ 0.00 8192
Wine 0.64 $\pm$ 0.01 0.62 $\pm$ 0.04 0.64 $\pm$ 0.04 0.63 $\pm$ 0.04 0.63 $\pm$ 0.04 0.63 $\pm$ 0.04 1599
Concrete 5.67 $\pm$ 0.09 5.23 $\pm$ 0.53 6.03 $\pm$ 0.58 5.28 $\pm$ 0.58 6.02 $\pm$ 0.65 5.47 $\pm$ 0.67 1030
Energy 1.80 $\pm$ 0.05 1.66 $\pm$ 0.19 2.09 $\pm$ 0.29 1.64 $\pm$ 0.36 2.53 $\pm$ 0.99 1.35 $\pm$ 0.97 768
Boston 3.01 $\pm$ 0.18 2.97 $\pm$ 0.85 3.28 $\pm$ 1.00 2.85 $\pm$ 0.72 3.54 $\pm$ 1.16 3.05 $\pm$ 0.91 506
Yacht 1.02 $\pm$ 0.05 1.11 $\pm$ 0.38 1.58 $\pm$ 0.48 2.52 $\pm$ 1.04 3.28 $\pm$ 1.39 1.62 $\pm$ 0.53 308
Comparison of aMCL and MCL. We can observe that aMCL performs comparably or outperforms vanilla MCL in most settings, especially for large dataset sizes, both in terms of distortion and RMSE. This outcome supports the claims made in the paper and is especially promising, given that the temperature scheduler was not specifically optimized for each dataset.
Comparison of aMCL and standard UCI baselines. We observe that aMCL performs on par with, and in some cases exceeds, standard baselines on the RMSE metric. This is noteworthy, as aMCL is not explicitly optimized for RMSE during training—its primary focus is on quantization. While perfect quantization would naturally result in optimal RMSE, achieving low RMSE is not the main objective of aMCL. For example, the RMSE performance of MCL is slightly worse across most datasets in those experiments.
Comparison with Relaxed-MCL. Finally, we compare aMCL and Relaxed-MCL, since aMCL can be interpreted as an adaptative version of Relaxed-MCL. Our results indicate that aMCL generally outperforms Relaxed-MCL in terms of distortion across nearly all datasets. However, we also observe that Relaxed-MCL demonstrates strong performance in terms of RMSE. We attribute these findings to the bias of Relaxed-MCL toward the conditional barycenter of the target distribution, which seems to improve RMSE at the expense of increased distortion. This trade-off arises because RMSE evaluates the accuracy of the barycenter estimation, while distortion measures the quantization performance. The trade-off between distortion and RMSE can be adjusted by tuning the value of $\varepsilon$. Further analysis of this parameter is presented in Appendix D.1, where we show that aMCL strikes a good balance between these two metrics.
These results strongly support the use of aMCL as a quantization algorithm in practical settings. To further evaluate its effectiveness, we also apply aMCL to a more challenging task, namely speech separation.
@@ 5.2 Application to speech separation
Speech separation consists of isolating each speaker’s signal from a mixture in which they are simultaneously active. This task is of major interest for automatic speech processing applications [50, 44, 71]. In these experiments, we explore the application of MCL and the proposed aMCL to the task of speech separation. An extensive description of the experiments is proposed in Appendix E.2.
@@ 5.2.1 Experimental setting
Dataset. Source separation experiments are conducted on the Wall Street Journal dataset [30] (WSJ0-mix), a standard benchmark for speech separation. We focus on the 2- and 3- speaker mixture scenarios, with each scenario including 20,000 training, 5,000 validation, and 3,000 testing mixtures.
Model architecture. The source separation task is solved using the Dual-Path Recurrent Neural Network (DPRNN) architecture [46]. DPRNN has been extensively used in speech separation, as it strikes a good balance between performance and number of trainable parameters [73, 65].
Separation metrics. We use the Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) to measure the separation quality [70, 39]. There is an ambiguity in finding the best assignment between predicted and active sources. The PIT SI-SDR loss [37] initially addresses this issue. We propose to use MCL and our new variant aMCL to perform this matching (See Appendix E).
@@ 5.2.2 Results
Table 3: 2- and 3- speaker source separation with PIT (topline), MCL, aMCL and Relaxed-WTA ($\varepsilon=0.05$). PIT SI-SDR metric ()($\uparrow$)( ↑ ) on the WSJ0-mix eval set. Results over three training seeds, with mean and standard deviation reported.
@@
Comparing PIT, MCL, Relaxed-WTA and aMCL. Table 3 presents the source separation results in the 2- and 3-speaker scenarios. First, aMCL demonstrates performance equivalent to or better than MCL. Both methods can be used for the separation task. However, we observed that MCL is subject to hypothesis collapse for some training seeds, while aMCL is more robust to initialization. This translates into a lower inter-seed standard deviation for aMCL. Second, we observe the advantage of aMCL over Relaxed-WTA, which is consistent with our previous analysis of the barycenter bias of this method. Third, aMCL performs equivalently to PIT in both scenarios. Note that by using MCL or aMCL, the number of predictions $n$ could exceed the number of sources $m$. This could be leveraged to improve the separation metrics (cf. Appendix E.3.2). Finally, aMCL improves the algorithmic complexity of PIT from $\mathcal{O}(m^{3})$ to $\mathcal{O}(mn)$ (cf. Appendix E.3.1). These results make aMCL stand as a good alternative to PIT.
Observing phase transition. When the metric is the Euclidean distance, the theoretical analysis of Section 4.3 and the synthetic experiments (see Figure 3) have highlighted a phenomenon of phase transition for aMCL. Here, we analyze the validation loss trajectory as a function of the temperature for different initial temperatures, and using an exponential scheduler. The curves with the two higher initial temperatures in Figure 8 exhibit a plateau until a given temperature. After this critical point, the loss decreases. Although the SI-SDR is non-Euclidean, this behavior resembles that observed for the Euclidean metric. This is detailed in Section E.3.3 of the Appendix.
@@ 6 Conclusion
This article introduces aMCL, a novel training method that combines deterministic annealing and the Winner-Takes-all training scheme to address two key issues of MCL: hypothesis collapse and convergence toward a suboptimal local minimum of its quantization objective. We provide a detailed analysis of aMCL’s training dynamics. Moreover, drawing on statistical physics and information theory, we provide insights into the trajectory of the aMCL predictions during training. In particular, we exhibit a phase transition phenomenon and establish its connection to the rate-distortion curve. We validate our analysis with experiments on synthetic data, on the UCI datasets, and on a real-world speech separation task. This demonstrates that aMCL is a theoretically grounded alternative to MCL in diverse settings. Future work includes a detailed analysis of the temperature schedule’s impact, the derivation of performance bounds, a thorough examination of our algorithm’s convergence, particularly at finite temperature, as well as an evaluation of its generalization capabilities on out-of-distribution samples.
Limitations. First, aMCL introduces a temperature schedule: this is a challenging hyperparameter tightly linked to the optimizer and its learning rate. The derivation of optimal schedules is left to future work. Second, annealing requires a slow temperature schedule to maintain model performance. This potentially leads to longer training times.
@@
7 Acknowledgments
This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854) and Hi! PARIS through their PhD in AI funding programs, and was performed using HPC resources from GENCI–IDRIS (Grant 2021-AD011013406R1). We are grateful to the reviewers for their insightful comments.
@@ A.1 Notations and motivations
Following the main paper notations, let $\mathcal{X}\subset\mathbb{R}^{d_{1}}$ and $\mathcal{Y}\subset\mathbb{R}^{d_{2}}$ denote the input and target spaces respectively. We will assume that $\mathcal{X}$ and $\mathcal{Y}$ are bounded. We note $p(x,y)$ the continuous density of a joint data distribution on $\mathcal{X}\times\ \mathcal{Y}$. Let $f=(f_{1},\dots,f_{n})\in\mathcal{F}_{\Theta}\subset\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})$ denote the hypothesis models, which are families of $n$ neural networks with parameters in $\Theta$. Likewise, let $\gamma=(\gamma_{1},\dots,\gamma_{n})\in\mathcal{F}_{\Theta^{\prime}}\subset\mathcal{F}(\mathcal{X},[0,1]^{n})$ denote the score models, with parameters in $\Theta^{\prime}$. We will sometimes write $\mathcal{F}=\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})$.
We denote by $\Delta_{n}$ the set of discrete distributions on $n$ items conditioned by points on $\mathcal{X}\times\mathcal{Y}$ (typically representing a soft assignment of a target to the hypotheses). We denote by $H(q)$ the entropy of a distribution $q$. If $q\in\Delta_{n}$, we write $H(q)=-\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}q(f_{k}|x,y)\log q(f_{k}|x,y)p(x,y)\mathrm{d}x\mathrm{d}y$. If $Z$ is a random variable with density $p(z)$ and $Y$ is another random variable, we define the differential entropy of $Z$ as $H(Z)=-\mathbb{E}_{Z}[\mathrm{log}(p(Z))]$ and its conditional entropy as $H(Z\mid Y)=-\mathbb{E}_{(Z,Y)}[\mathrm{log}(p(Z\mid Y))]$. Mutual information $I(Z,Y)$ between two random variables $Z$ and $Y$ is defined as $I(Z;Y)=H(Z)-H(Z\mid Y)$.
Let $t\in\mathbb{N}$ denote a training step, and $t_{\mathrm{epoch}}\in\mathbb{N}\cup\{\infty\}$ denote the number of training epochs. Let $T:t\mapsto T(t)\geq 0$ be a temperature schedule that decreases with the training step and verifies $\mathrm{lim}_{t\rightarrow t_{\mathrm{epoch}}}T(t)=0$. Note that we consider the temperature to be constant across an epoch, with linear or exponential decrease between epochs. Let $\ell:\mathcal{Y}^{2}\rightarrow\mathbb{R}_{+}$ denote an underlying loss function. We will restrict our analysis to the Euclidean squared distance $\ell(\hat{y},y)=\lVert y-\hat{y}\rVert^{2}$ unless otherwise stated. For any vector $z$, we note $z^{t}$ as its vector transposition. Gradient descent involves a learning rate schedule, that we note $\lambda_{t}$.
@@
which can be seen as a generalization of the conditional distortion [56]. The two definitions coincide when $p(y\mid x)$ is independent of $x$. Minimizing (16) is NP-hard [1, 10, 48]. The K-means algorithm, which is the standard approach for this task, can take an exponential number of steps before converging [69], and is only guaranteed to find a local minimum in general settings [13, 7]. The same applies to WTA: due to its greedy nature, it can be trapped, at any point of its training trajectory, into a local minimum that it will never manage to escape. One common issue is the hypothesis collapse [57, 18], a situation in which some of the hypotheses are left unused, therefore leading to a suboptimal configuration (see Figure 1).